Find the velocity of a transverse wave in a string which is 4.0 m long with a mass with a total mass of 20 g when under a tension of 100N


1
Expert's answer
2021-12-07T09:54:19-0500
"v=\sqrt{\frac{100(4)}{0.02}}=141\frac{m}{s}"